In a two good economy, only burritos and t-shirts are sold. In 2014, the base year, the price of a burrito was $7.25 and 10 burritos were sold. Meanwhile, in 2014, the price of a t-shirt was $5.00, and 100 were sold. In 2015, the price of a burrito was $8.50 and 8 burritos were sold, and the price of a t-shirt was $4.00 and 110 t-shirts were sold. What was the GDP for 2014? What was the nominal GDP (NGDP) for 2015 and the real GDP (RGDP) for 2015?
Answer #1

Answer

Real GDP =sum of (quantity of the year * price of base year)

Nominal GDP=sum of (quantity of the year * price of the year

the base year is 2014 so the nominal and real GDP is the same for the year

GDP in 2014 =7.25*10+5*100=$572.5

Nominal GDP in 2015 =8.5*8+4*110=508

real GDP in 2015 =7.25*8+5*110=608